Solve for x in the triangle: 612 3 12 N 6 2V6 24
Answer #1

Solution Given L A245, LC=30 find = x LA+LB TCC = 180 (sume of all interior 45+30+ LB = 180 angles are 180.) a LB = 180-75 a=